# HyResPINNs: Adaptive Hybrid Residual Networks for Learning Optimal Combinations of Neural and RBF Components for Physics-Informed Modeling

### v1

@@ 3.1 Hybrid Residual Blocks
where ${\bf x}^{c}_{i}$ is the center of the $i$-th RBF, $\tau$ is a scaling parameter, and $\|{\bf x}-{\bf x}^{c}_{i}\|$ represents the Euclidean distance between ${\bf x}$ and ${\bf x}^{c}_{i}$. The Wendland kernel is compactly supported, meaning $\phi_{i}(\mathbf{x})=0$ for $\|{\bf x}-{\bf x}^{c}_{i}\|\geq\tau_{i}$, which leads to sparse kernel matrices and computational efficiency. Each $\tau_{i}$ is a trainable parameter in the network and optimize through gradient descent along with all other network parameters. Figure 3 demonstrates learned scale parameters within an optimized RBF kernel.
Block Neural Network
Following the convention of Cyr et al. (2020), we represent the output of the DNN $F_{N}^{(l)}({\bf x}^{(l)})$ $F_{N}^{(l)}\in\mathbb{R}^{d}\rightarrow\mathbb{R}$ of width $w$, as a linear combination of adaptive basis functions given by
$F_{N}^{(l)}({\bf x};a,\theta^{h})=\sum_{i=1}^{w}a_{i}\sigma_{i}{\left({\bf x};\theta^{h}\right)},$ (14)
where each $a_{j}$ for $j=1,..,w$ and $\theta^{h}$ constitute the weights and biases in the last layer and hidden layers respectively, forming the set of all network parameters $\theta$. Then, each $\sigma_{j}$ are non-linear smooth activation functions such as Tanh acting on the outputs of the hidden layers. We choose to use a standard DNN architecture, but any DNN architecture such as ResNets, would work. The parameters $\theta$ are computed through some iterative optimization technique. In this work, we use variants of gradient descent methods such as ADAM Kingma & Ba (2015) and L-BFGS Liu & Nocedal (1989).
Figure 2 shows a visualization of our full model architecture with adaptive residual block skip connections, along with the input block structure used to lift the inputs to the desired higher dimension, and the output neural network block used to project each block’s output down to the output dimension.
@@
Table 1: Relative $L^{2}$ test error results for various PDE problems and baseline methods.
@@ 4 Experimental Results and Discussion
In this section, we demonstrate the effectiveness of the proposed HyResPINNs  architecture across a diverse collection of benchmark problems compared to the leading baseline methods. We show that HyResPINNs  consistently produce lower relative $L^{2}$ errors between the predicted and ground truth solutions for the same training set size when compared to the applicable baselines trained under the same experimental procedures and hyperparameters. We compare each method on the 1D non-linear hyperbolic Allen-Cahn equation and the Darcy Flow equation in two and three dimensions with both Dirichlet and Neumann boundary conditions. Our main results are summarized in Table 1.
@@
Figure 4: 2D Allen-Cahn equation: Comparison of the predicted solutions for the Allen-Cahn equation using HyResPINN, standard PINN, and RBF network models. The top row shows the predicted solutions at for HyResPINN (left), standard PINN (center), and RBF network (right). The second row illustrates the absolute error between the predicted solutions and the exact solution. The bottom row shows the predicted solutions for three time steps comparing the exact solution (green), HyResPINN (orange), and standard PINN (dashed blue). The HyResPINN model effectively captures both the smooth and sharp features of the solution, while the standard PINN struggles with sharp transitions, leading to larger errors in these regions.
@@
Baseline Methods. We compare the proposed HyResPINN  against a suite of baseline methods under exactly the same hyper-parameter settings in each example. Baselines include the standard PINN as originally formulated in (Raissi et al., 2019) (PINN), a PINN based on the architectural guidelines proposed in Wang et al. (2023) (ExpertPINNs), PINNs with residual connections (ResPINNs), PirateNets Wang et al. (2024a), and the stacked PINN approach of (Howard et al., 2023) (StackedPINNs). To ensure a fair comparison, we implemented the approaches detailed in (Raissi et al., 2019; Wang et al., 2024a; 2023; Howard et al., 2023) following the architectural details provided in each.
@@
Experimental Setup. We follow similar experimental design procedures as those described in Wang et al. (2022a; 2023; 2024a). A full description of the experimental setups are included in Appendix A, and exact hyper-parameter setting are detailed in Table 2. We train all models on Nvidia A100 GPU running Centos 7.2. Code for our methods and all compared baseline approaches is written in libtorch $1.11$ (the C++ version of PyTorch) and will be made publicly available upon publication.
4.1 1D Allen-Cahn
We first focus on the Allen-Cahn equation, a challenging benchmark for conventional PINN models that has been extensively studied in recent literature  (Wight & Zhao, 2020; Wang et al., 2022a; Daw et al., ). For simplicity, we consider the one-dimensional case with a periodic boundary condition with $t\in[0,1]$, and $x\in[-1,1]$:
@@
The Allen-Cahn PDE is an interesting benchmark for PINNs as it introduces periodic boundary conditions, and because it is a “stiff” PDE that challenges PINNs to approximate solutions with sharp space and time transitions.
Hybrid residual blocks improve prediction accuracy.
We first demonstrate that the hybrid residual block structure introduced in our method enhances the ability to capture both smooth and sharp features in the 2D Allen-Cahn solution. As shown in Figure 4, the proposed approach outperforms competing baselines, which struggle to accurately represent the sharp transitions in the solution. This difficulty arises from the smoothness constraints of standard neural network architectures. In contrast, our method effectively captures these sharp transitions by leveraging a combination of smooth neural networks and adaptive non-smooth RBF kernels, allowing the model to balance between global smooth behavior and localized discontinuities. Figure 3 shows a random subset of the learned RBF kernels captured by the proposed hybrid approach. Specifically, it captures similar RBF structures (to a standard RBF network), but benefits from the additional flexibility provided by the NN component, allowing for more nuanced function approximation in regions requiring a balance of smoothness and sharp transitions. This is likely apparent in the on average smaller smaller kernels learned by the hybrid block.
Stacking structure of residual blocks improve prediction accuracy.
We next show that the stacking structure employed by the HyResPINN  in each hybrid residual block structure introduced here further helps to capture the smooth and sharp features in the Allen-Cahn solution. Figure 5 shows the predicted solutions between the proposed method and the competing baselines. The competing methods need help accurately capture the sharp transitions in the solutions. In contrast, our approach more accurately captures these sharp transitions due to the conjunction of the smooth neural network architecture and the learned non-smooth RBF kernels.
Figure 6: Allen-Cahn equation: Comparison of mean relative $L^{2}$ error across various methods, plotted against the number of training iterations (left) and the mean training time (right). Vertical bars in right plot indicate the time for which each method achieved an error of $210^{-}$ 2.
We next evaluate the effect of network depth on predictive accuracy for different architectures. Figure 5 HyResPINN consistently achieves the lowest error across different network depths. The performance of other methods shows varying sensitivity to the number of hidden layers, with HyResPINN offering the most robust performance. Further, the left plot in Figure 6 shows that HyResPINN and PiratePINN achieve the lowest error over iterations, with HyResPINN consistently reducing the error more effectively than the other methods. The right plot demonstrates that—although HyResPINN has a slightly higher training cost—it significantly outperforms other methods in accuracy, particularly for long training times. These results highlight the robustness and efficiency of HyResPINN in solving complex PDEs.
@@
While attempting to replicate the results of (Wang et al., 2024a), using the reported hyperparameters and implementation details, we observed a slight performance gap between the results obtained in our implementation and those listed in their paper. The observed discrepancy in the final model’s performance could stem from several factors such as subtle variations in the implementation environment—such as hardware differences, software library versions, or random seed initialization. Despite the differences between our replication results and those reported in (Wang et al., 2024a), the comparison between our architecture and the replicated version of their approach remains fair, as both were tested under identical conditions using the same dataset, hyperparameters, and training setup (including hardware and software environment). Our findings indicate that our architecture outperforms the replicated version of their model in many cases, along with all other replicated baseline methods.
4.2 Darcy Flow

### v2

@@ 3.1 Hybrid Residual Block Structure
where $\mathbf{x}_{i}^{c}$ represents the center of the $i$-th RBF, $\tau_{i}$ is a trainable scaling parameter, and $r_{i}=\|\mathbf{x}-\mathbf{x}_{i}^{c}\|$ corresponds to the Euclidean distance between $\mathbf{x}$ and $\mathbf{x}_{i}^{c}$. Each $\tau_{i}$ is optimized through gradient descent, allowing the network to dynamically adapt the kernel shape throughout training. Figure 3 demonstrates a subset of the learned scale parameters for an optimized RBF kernel.
Figure 3: 2D Allen-Cahn equation: Comparison of two random RBF kernels from a standard RBF-Net (top row) and the first block of a HyResPINNs (bottom row). Each subplot shows the (final) learned RBF scale parameter ($\tau$) for the selected center points (marked by red crosses).
NN Block Component.
The NN component is defined as:
$F_{N}^{(l)}(\mathbf{x})=\mathcal{NN}(\mathbf{x};\theta),$ (10)
where $\theta$ represents the set of trainable parameters in the network. The NN consists of at least one hidden layer with a non-linear activation function, which enables the model to approximate complex functions while maintaining smoothness. Unlike the RBF component, which specializes in localized features, the NN component is designed to capture global trends in the solution. Although a standard deep neural network is used in this work, other architectures such as residual networks (ResNets) could be employed in place of the NN component within the hybrid residual block.
Adaptive Block Parameter.
To balance the contributions of $F_{R}^{(l)}$ and $F_{N}^{(l)}$, we employ a trainable weighting function $\phi(\alpha^{(l)})$, where $\phi$ is the sigmoid function, ensuring that $\phi(\alpha^{(l)})$ remains within the range $[0,1]$. We initialize $\phi(\alpha^{(l)})=0.5$, ensuring that both components contribute equally at the start of training. The learnable parameter $\alpha^{(l)}$ dynamically adjusts throughout training to optimize the relative influence of the NN and RBF components based on the characteristics of the problem.
3.2 Deep Hybrid Residual Architecture
The full HyResPINN architecture, shown in Figure 2, consists of multiple hybrid residual blocks stacked within a residual learning framework. Each block’s output is propagated forward, with trainable skip connections modulated by $\phi(\beta^{(l)})$. The model computes the intermediate representations $\bar{\mathbf{a}}^{(l)}$ at each layer according to:
$\bar{\mathbf{a}}^{(l)}=\phi(\beta^{(l)})F^{(l)}(\mathbf{x})+(1-\phi(\beta^{(l)}))\bar{\mathbf{a}}^{(l-1)}.$ (11)
After passing through $L$ hybrid residual blocks, the final prediction is obtained via a deep neural network (DNN), which refines the hybrid representation into the final output:
$\mathbf{y}=\text{DNN}(\bar{\mathbf{a}}^{(L)}).$ (12)
Unlike PirateNets (Wang et al., 2024), which initialize their residual connection parameter to zero at the start of training, HyResPINN initializes each $\phi(\beta^{(l)})=1$. This initialization ensures that the early training dynamics do not force the network to learn non-linearities from scratch, but rather begin with a meaningful blend of residual and block outputs. PirateNets also use an additive residual formulation rather than the convex combination approach in HyResPINN, which allows our method to explicitly control the contribution of each block’s output through a learned weighting function. This structured balancing mechanism provides greater flexibility in adapting to both smooth and non-smooth solution structures.
Regularization.
To encourage smoothness and prevent excessive non-linearity, we introduce an $L_{2}$ penalty on the trainable parameters $\alpha$ and $\beta$:
$\mathcal{L}$ $=\lambda_{i}\mathcal{L}_{i}+\lambda_{b}\mathcal{L}_{b}+\lambda_{r}\mathcal{L}_{r}+\lambda_{p}\sum_{i=1}^{N_{blocks}}\left(\alpha_{i}^{2}+\beta_{i}^{2}\right).$ (13)
This regularization term helps stabilize the training process by discouraging excessive reliance on either the NN or RBF component in each block. By contrast, PirateNets do not apply explicit regularization to their trainable residual parameters, which can lead to less controlled adaptation between blocks. The additional regularization in HyResPINN ensures that the learned weighting functions $\phi(\alpha^{(l)})$ and $\phi(\beta^{(l)})$ remain well-behaved throughout training, preventing overfitting to sharp local variations and promoting a more balanced function approximation.
@@ 3.2 Deep Hybrid Residual Architecture
Table 1: Overview of the relative $L^{2}$ test error results for various PDE problems and baseline methods.
@@ 4 Experimental Results and Discussion
In this section, we demonstrate the effectiveness of the proposed HyResPINNs  architecture across a diverse collection of benchmark problems compared to the leading baseline methods. We first tested each method on the 1D non-linear hyperbolic Allen-Cahn equation to examine how the local RBFs in the HyResPINN  blocks help to predict sharp solution transitions. We then tested the Darcy Flow equation in 2D and 3D annulus/extruded annulus domains (respectively) with Neumann and Dirichlet boundary conditions to assess how HyResPINNs predict solutions on irregular domains, evaluating its robustness in handling geometric complexities. Next, we examined the 2D Darcy Flow problem with rough coefficients to evaluate how our method handles discontinuities and non-smooth solutions. Finally, we tested the Kuramoto–Sivashinsky equation to test our method on regular and chaotic problems, demonstrating its capability to model complex spatiotemporal dynamics. Appendix B details all PDE benchmarks and corresponding experimental settings. Table 1 summarizes our main results.
@@
Figure 4: Allen-Cahn equation: Comparison of the predicted solutions for the Allen-Cahn equation using HyResPINN, standard PINN, and RBF network models. The top row shows the predicted solutions for HyResPINN (left), standard PINN (center), and RBF PINN (right). The second row shows the absolute error between the predicted and true solutions. The bottom row shows the predicted solutions for time steps ($t=0.25,0.5,0.99$) compared to the true solution.
@@
Baseline Methods. We compare the proposed HyResPINN  against a suite of baseline methods, including 1) the standard PINN as originally formulated in (Raissi et al., 2019b) (PINN), 2) a PINN based on the architectural guidelines proposed in Wang et al. (2023) (ExpertPINNs), 3) PINNs with residual connections (ResPINNs), 4) PirateNets Wang et al. (2024), and 5) the stacked PINN approach of (Howard et al., 2023) (StackedPINNs). To ensure a fair comparison, we implemented the approaches detailed in (Raissi et al., 2019b; Wang et al., 2024; 2023; Howard et al., 2023) following the architectural details provided in each.
@@
Experimental Setup. We follow similar experimental design procedures as those described in Wang et al. (2022a; 2023; 2024). The code for our methods and all compared baseline approaches is written in LibTorch $1.11$ (the C++ version of PyTorch) and will be publicly available upon publication. We train all models on a Nvidia A100 GPU running Centos 7.2. A full description of the experimental setups are included in Appendix C.
4.1 Allen-Cahn
First, we present results for 1D the Allen-Cahn problem with periodic boundary conditions, a well-known challenge for conventional PINN models and a widely studied benchmark in recent literature  (Wight & Zhao, 2020; Wang et al., 2022a; Daw et al., ). This problem is particularly difficult for two reasons. First, the Allen-Cahn PDE requires PINNs to approximate solutions with sharp transitions in space and time, posing challenges for standard neural network architectures. Second, there is an inherent incompatibility between the initial and boundary conditions, leading to a weak discontinuity at time $t=0$. Traditional numerical methods would likely produce oscillations near the discontinuity due to the mismatch between the initial and boundary conditions. However, unlike traditional solvers, PINNs do not rely on discretization and instead learn their model parameters through optimization. This optimization process may mitigate spurious oscillations, as it relies on stochastic updates and continuous function approximations. Specifically, implicit regularization from both mini-batch training and stochastic gradient descent (SGD) may contribute to stabilizing optimization and reducing any spurious oscillations Sekhari et al. (2021). Therefore, despite this discontinuity, we evaluate our method on this benchmark due to its established role in assessing PINN performance.
Hybrid Residual Blocks Improve Prediction Accuracy.
Figure 4 demonstrates the advantages of the hybrid residual block structure in capturing both smooth and sharp features in the 1D Allen-Cahn solution. The absolute error plots show that HyResPINN exhibits significantly lower errors than standard PINNs and RBF-PINNs, validating its improved accuracy. As expected, the PINN prediction provides a smooth approximation but struggles with sharp transitions due to the inherent smoothness of neural networks. Meanwhile, RBF-Net fails to recover both the global structure and fine local details because, while its length scale is learned through optimization, its center locations and number of basis functions remain fixed. This rigid structure prevents it from dynamically placing basis functions in regions of rapid change, leading to poor approximation of sharp features.
HyResPINN resolves these limitations by combining a smooth neural network with localized RBF adaptivity. The neural network component ensures global consistency, while the RBF kernels allow for finer, more flexible local approximations. This synergy enables the model to balance smooth and non-smooth behavior, significantly improving predictive accuracy.
Figure 6: Allen-Cahn equation: Comparison of mean relative $L^{2}$ error across various methods, plotted against the number of training iterations (left) and the mean training time (right). Vertical bars in right plot indicate the time for which each method achieved an error of $10^{-2}$.
Depth in Residual Blocks Improves Accuracy.
Figure 5 demonstrates that increasing the number of residual blocks in HyResPINN leads to consistently lower errors, outperforming the competing methods. Residual learning allows deeper networks to refine predictions iteratively, progressively improving both global structure and sharp transitions. The competing methods struggle to maintain accuracy as depth increases, whereas HyResPINN remains robust, demonstrating the effectiveness of deep hybrid residual learning.
Training Efficiency and Convergence of HyResPINNs.
Figure 6 (left) shows that HyResPINN achieve the lowest error over iterations, demonstrating the most consistent reduction in error. This suggests that the hybrid residual structure efficiently utilizes depth to refine solutions progressively.
While the right plot in Figure 6 shows that HyResPINN has a slightly higher training cost than other methods, its significantly lower error justifies this cost. In particular, for long training times, HyResPINN consistently outperforms all baselines, highlighting its robustness in solving complex PDEs. The vertical bars indicate the training times required for each method to reach an error threshold of $10^{-2}$, further demonstrating HyResPINN’s efficiency in achieving high-accuracy solutions. This makes the training cost vs. accuracy trade-off more explicit and clearly separates training efficiency from accuracy discussions.
4.2 Darcy Flow with Smooth Coefficients
We next test the performance of HyResPINN on the Darcy Flow equation with smooth coefficients. Darcy Flow serves as a fundamental model for various physical processes, including porous media flow, heat transfer, and semiconductor doping. Since real-world applications often involve solving these equations under varying boundary conditions, it is essential to evaluate both solution accuracy and robustness to different constraints. To assess convergence, we construct smooth manufactured solutions in two settings: a 2D annular domain and a 3D cylindrical domain obtained by extruding the 2D geometry along the z-axis. The exact solutions are given as,
$u(x,y)$ $=\sin(x)\sin(y),$ (14)
$u(x,y,z)$ $=\sin(x)\sin(y)\sin(z),$ (15)
for the 2D and 3D cases, respectively. By substituting these exact solutions into the model problem we define the source term and the boundary data. The true solutions are shown in Figure 9 for both the 2D and 3D cases. The detailed problem setup, including boundary conditions and parameter choices, is outlined in Appendix B.2.
We perform a convergence study by analyzing the mean relative $L^{2}$ error as the number of training collocation points increases. The problem is solved under both Dirichlet and Neumann boundary conditions to test robustness. Figure 7 presents the results for both 2D and 3D cases, where solid lines represent the solution error, and dashed lines correspond to the x-directional flux errors.
This experiment allows us to evaluate not only the accuracy of HyResPINN but also its ability to efficiently utilize training data. A robust method should achieve lower errors with fewer training points while remaining stable across different boundary conditions.
The results in Figure 7 demonstrate that HyResPINN consistently achieves the lowest errors compared to baseline methods across all training sizes. Notably: HyResPINN achieves high accuracy with fewer training points, demonstrating its ability to generalize well with limited data. Unlike traditional PINN-based approaches that may exhibit sensitivity to Dirichlet or Neumann constraints, HyResPINN remains stable across different boundary types. The method extends seamlessly from 2D to 3D without significant degradation in performance, confirming its suitability for large-scale PDE problems. These results further validate the effectiveness of hybrid residual learning in solving elliptic PDEs with smooth solutions.
4.3 Darcy Flow with Rough Coefficients
A key challenge when solving the Darcy Flow equation 16 in scenarios with discontinuous coefficients is accurately approximating the flux variable, particularly its normal and tangential components. The normal component of the flux remains continuous across material interfaces, whereas the tangential component may exhibit discontinuities. This behavior is crucial for physically accurate solutions, especially when solving problems involving heterogeneous materials or rough coefficients.
In contrast, traditional collocated methods—such as Galerkin, stabilized Galerkin, and least-squares finite elements—often struggle to capture these discontinuities in flux, typically resulting in oscillations at material interfaces. These oscillations indicate that the methods fail to accurately respect the physical continuity of the flux field.
In this section, we demonstrate that HyResPINN and its variants effectively capture the physical behavior of flux, providing accurate approximations of both the normal and tangential components in the presence of discontinuous coefficients. This is particularly evident in the convergence analysis for the "five strip problem", a well-documented manufactured solution test Nakshatrala et al. (2006); Masud & Hughes (2002), which evaluates the ability of a method to preserve the continuity of normal flux across material interfaces. As shown in Figure 10, the true solutions for both the solution field and the flux components are depicted for reference. We solve the problem with Neumann boundary conditions, which are commonly used for this problem setup.
Convergence results are presented in Figure 8, where we compare HyResPINN against baseline methods. The results show that HyResPINN is capable of accurately capturing the solution and flux, though it requires larger training sizes compared to the other methods. Nonetheless, it outperforms the baselines in terms of accuracy and physical fidelity, particularly in approximating the flux field without introducing undesirable oscillations.
4.4 Kuramoto–Sivashinsky Equation
In this section, we use the Kuramoto–Sivashinsky equation as a benchmark problem under periodic boundary conditions. The detailed problem setup is provided in Appendix B. The Kuramoto–Sivashinsky equation is a fourth-order nonlinear partial differential equation widely used to model spatiotemporal patterns in fluid dynamics, combustion, and other systems exhibiting chaotic behavior. Solving this equation allows us to assess HyResPINN’s ability to approximate highly complex and chaotic systems. HyResPINN achieved relative $L_{2}$ errors of $15.610\%$ for the regular case and $64.50\%$ for the chaotic case, while ExpertPINNs achieved significantly lower errors of $2.120\%$ and $3.272\%$, respectively. These results highlight the need for further refinement of HyResPINN, particularly in handling chaotic regimes.
5 Conclusion
In this work, we introduced HyResPINNs, a novel class of physics-informed neural networks that incorporate adaptive hybrid residual blocks combining the strengths of standard neural networks and radial basis function (RBF) networks. Our architecture effectively leverages smooth and non-smooth function approximators. The adaptive combination parameters within each block allow the model to balance the contributions of neural and RBF components during training. Our experiments demonstrate that HyResPINNs  outperform traditional PINNs and state-of-the-art methods in accuracy and robustness, particularly for problems involving mixed smooth and non-smooth regions. This work increases the flexibility and generalizability of PINNs by bridging the gap between classical numerical methods and DNN-based approaches for solving PDEs.
Appendix A Computational Cost
Allen-Cahn 1D Space/Time Periodic $34.99$ $17.65$ $64.81$ $22.29$ $116.72$ $150.88$
DarcyFlow 2D Annulus Neumann $2.37$ $2.76$ $18.67$ $2.18$ $11.88$ $5.73$
(smooth coefficients) Dirichlet $2.08$ $2.09$ $17.64$ $1.64$ $10.87$ $11.87$
3D Annulus Neumann $8.1$ $8.5$ $27.6$ $6.9$ $20.8$ $35.9$
Dirichlet $5.9$ $6.0$ $21.3$ $5.2$ $15.3$ $30.1$
(rough coefficients) 2D Box Neumann $2.4$ $2.4$ $7.7$ $2.1$ $7.7$ $8.4$
Table 2: Average wall-clock training time (in minutes) for each PDE problem and baseline methods.
Appendix B Benchmark PDEs
B.1 Allen-Cahn Equation
For simplicity, we consider the one-dimensional case with a periodic boundary condition with $t\in[0,1]$, and $x\in[-1,1]$:
@@ B.1 Allen-Cahn Equation
While attempting to replicate the results of (Wang et al., 2024), using the reported hyperparameters and implementation details, we observed a slight performance gap between the results obtained in our implementation and those listed in their paper. The observed discrepancy in the final model’s performance could stem from several factors such as subtle variations in the implementation environment—such as hardware differences, software library versions, or random seed initialization. Despite the differences between our replication results and those reported in (Wang et al., 2024), the comparison between our architecture and the replicated version of their approach remains fair, as both were tested under identical conditions using the same dataset, hyperparameters, and training setup (including hardware and software environment). Our findings indicate that our architecture outperforms the replicated version of their model in many cases, along with all other replicated baseline methods.
B.2 Darcy-Flow Equation
In the Allen-Cahn experiment, we employed mini-batching with randomly sampled collocation points. However, for the Darcy Flow experiments, we aimed to systematically study the effect of increasing the number of collocation points on convergence. To achieve this, we used static training datasets, ensuring that all experiments maintained a fixed set of training points per run. This allows for a controlled evaluation of error reduction as the dataset size increased, as shown in Figure 7. The use of static datasets also eliminated variability from stochastic sampling, enabling clearer comparisons between problem dimensions and boundary conditions. This approach provides a more precise understanding of the trade-off between accuracy and the number of training points, highlighting the efficiency of HyResPINN in solving elliptic PDEs.
B.2.1 Smooth Coefficients
Figure 9: Smooth Darcy Flow equation: True solutions of the 2D and 3D problems.